OpenStudy (anonymous):

Please Help! The sample space of an experiment is [A,B,C,D] and P(A)=0.1, P(B)=0.3, P(C)=0.4, P(D)= a.)0.3 b.)0.4 c.)0 d.)0.2

OpenStudy (anonymous):

All the probabilities in the sample space must add up to 1. \[P(A)+P(B)+P(C)+P(D)=1~~\implies~~P(D)=1-0.1-0.3-0.4=?\]